OpenStudy (anonymous):

(2x+3)(2x-3) how do i use the foil method

OpenStudy (amistre64):

the same way youd use the lfio method or 22 other methods :)

OpenStudy (anonymous):

Get the items in the first set of parenthesis and times the out one at a time and then total.

OpenStudy (amistre64):

just multiply the 2 together ...

OpenStudy (anonymous):

\[(2x)(2x)+(2x)(-3)+3(2x)+3(-3)\]

OpenStudy (anonymous):

4x^2-9

OpenStudy (aravindg):

x= -3/2 or 3/2

OpenStudy (amistre64):

im old and tend to get lost in the f the o the i and the l so I just do it the simplest way I know how :) 2x+3 2x-3 --------- -6x-9 4x^2+6x ----------- 4x^2 + 0 -9 4x^2 - 9

OpenStudy (anonymous):

i hate math ugh ....

OpenStudy (anonymous):

i love how all of a sudden x equals something!

OpenStudy (anonymous):

you might could recognize that \[(a+b)(a-b)=a^2-b^2\] foil or no foil. here you have \[(2x+3)(2x-3)=(2x)^2-3^2=4x^2-9\]

OpenStudy (anonymous):

Well at least you're reaching out for help and more importantly explanation. Most folks here just wants answers.

OpenStudy (anonymous):

well i don't wanna havehelp all my life i like to be independent

OpenStudy (anonymous):

I prefer to just use normal distribution: \[(2x+3)(2x-3)\]\[=2x(2x-3) + 3(2x-3)\]\[=4x^2 - 6x + 6x - 9\]\[=4x^2 - 9\]

OpenStudy (anonymous):

ok thanks for the help